Chapter I THESIS OBJECTIVES

2

Chapter I

THESIS OBJECTIVES

1. Motivation

I could begin by recounting all the scientifically compelling reasons why the merging process is so important, both locally for understanding the structure of our own Galaxy, and cosmologically for the implications it has for the evolution of structure in the universe. I could wax poetic about the wondrous ability of the Very Large Array (VLA) to gather a wealth of data over large areas in a reasonable amount of time, making it uniquely suited for studying these systems which throw large amounts of neutral hydrogen about with such apparent abandon. I could boast about how such data will prove to be a critical link between the observational and theoretical merger worlds. And I could lie and say that it was the realization that interactions had to be important that drove me to define this thesis.

But this is the motivation section. And what motivated me was not an appreciation of the role interactions play in the universe or a realization of the types of data that were needed to study them. What motivated me was the dumbfound wonderment I experienced after discovering Halton Arp’s “Atlas of Peculiar Galaxies” while an undergraduate at Case Western Reserve University, and the compelling beauty of the morphology and kinematics of the H i in the interacting system Mkn 348 as viewed in intensity-hue mode on an I^{2}S that I witnessed as a summer student at the VLA in 1985. And I knew that I wanted to make the most beautiful thesis I could ever imagine.

And I believe I can say, with perhaps an undue lack of modesty, that I have been privileged with the opportunities, encouragement, support, and collaborators to do just this. I will presently discuss the scientific goals and relevance of this work to the field. But I’ll be lucky if these words last through this century. The figures are my contribution to the field, and the desire to get them so I could stare again in dumbfound wonderment was my motivation.

2. Historical Perspectives

Much attention has been focused through the years, in various degrees and to varying extent, on the extragalactic mutations that fall under the broad yet appropriate moniker of “Peculiar Galaxies”. From the very earliest days of the Hale and Lick giant telescopes, these objects elicited awe, bewilderment, and fanciful musings from the observing astronomers. Galaxies in birth (Burbidge et al. 1963), galaxies in death, galaxies fragmenting (Voronstov-Vel’yaminov 1959), exploding or imploding (Zwicky 1970), ejecting (Arp 1966; Voronstov-Vel’yaminov 1959), or at the mercy of a complicated nest of electromagnetic fields and forces (Pikel’ner 1968); all have been proposed to explain the nature and origin of these systems.

The idea that interactions between galaxies are important has been around for some time. However, it took years of careful documentation for enough evidence to be gathered to bring this idea into the mainstream. It wasn’t long after the concept of galaxies as “Island Universes” was accepted that Holmberg realized that galaxies are not quite islands after all and performed his analog computation of the gravitational force between pairs of them (Holmberg 1941). Observationally, it was Zwicky’s extensive work (1950, 1956, 1964) and characteristic insight^{1} in referring to the tails and bridges that he found in his surveys as tidal, the attention paid to “young” looking galactic systems by the Burbidge’s et al. (1963 and references

therein), and the compilation of the “Atlas and Catalogue of Interacting Galaxies” by

Voronstov-Vel’yaminov (1959) and the “Atlas of Peculiar Galaxies” by Arp (1966), that gave astronomers the observational evidence needed to draw attention to these misfits of the Hubble sequence.

2.1 Bridge and Tail Building.

Aided by the growing power of digital computers, the needed attention was brought to bear in the 60’s and 70’s (see reviews in Schweizer 1990, Barnes & Hernquist 1992, Barnes 1994). These early computer algorithms simulated the effects of gravity on a set of massless test particles moving in the combined field of two rigid potentials. Due to its simplicity, this “restricted three body” model enabled many different encounter scenarios to be run in a relatively short time. It wasn’t until fully self-consistent simulations were run (van Albada & van Gorkom 1977, White 1978) that the extreme stickiness of galaxies was fully demonstrated (although it was estimated by both Holmberg in 1941 and Alladin in 1965). However, these simple codes were more than adequate to describe the early development of interacting systems and allowed some particularly insightful astronomers to anticipate results that would take over a decade to demonstrate numerically.

Toomre & Toomre’s seminal work (1972, hereafter referred to as TT) stands out among the others due to their appreciation of the need for slow, prograde orbits of bound pairs to produce a substantial tidal response. This work was also ground-breaking in that it departed from previous theoretical work and actually succeeded in reproducing the tidal morphology of four interacting systems.^{2} Since two of these systems figure prominently in the present work, we reproduce TTs final solutions here in figure 1 for Arp 295 and in figure 2 for NGC 4676 (their

1 More correctly, Directed Intuition, cf. Zwicky 1971.

2 The Toomres’ relatively easy success has probably contributed to the too prevalent notion that morphological peculiarity alone suffices as evidence for gravitational interaction. The Toomres themselves state: “We ourselves obviously suspect most such peculiarities to be the results of tides. Yet even to us that resemblance seems unmistakable in only about 30 of the 300-plus examples contained in Arp’s atlas” [their italics] (TT p. 648). So while a disturbed morphology may suggest a tidal origin, kinematic data are needed to fully justify such a claim.

FIGURE 1: Model of Arp 295 from restricted three-body simulation of Toomre & Toomre (1972, figure 19) from different viewing perspectives (β=0^{o} is face-on to the orbital plane, while β=90^{o} is edge-on). Particles were placed around only one system (corresponding to Arp 295a), with the secondary intruder represented by a single point source of one-quarter the mass of the primary. The view at the bottom is quite similar to the actual appearance of the Arp 295 system (Fig. 1a of Chapter II).

FIGURE 2: Model of NGC 4676 from restricted three-body simulation of Toomre & Toomre (1972, figure 22) from different viewing perspectives. The top view is face-on to the orbital plane, while the bottom is the best match to the optical morphology (Fig. 1b of Chapter II). This encounter was between two equal mass disks with inclinations of 15^{o} (NGC 4676a) and 60^{o} (NGC 4676b).

figures 19 and 22, respectively). In the Arp 295 simulation, particles were only placed in orbit around one of the galaxies (Arp 295a, the southern disk). In both figures, the percentages indicate the original radius of the particles in terms of the distance of closest approach, R_{min}. Lines connecting particles originally at 0.6R_{min} in each galaxy are drawn in Fig. 2.

This work included an extensive survey of the effect of varying orbital parameters and an exhaustive discussion of the process of bridge and tail making. Most of the conclusions hold true today, and I repeat here some of the particularly useful results to keep in mind when viewing such systems. The quotes in the following denote terms or phrases used specifically in TT.

• Because “simple gravity” is the mechanism for creating such features, the forces are symmetric, with the tails arising from material that was “left behind” on the “antitidal” side of the disk (that half of the disk which is on the far side of the companion at orbital periapse), and a bridge arising from particles which are “torn away” from the near side. This is illustrated here in figure 3, again taken from TT (their figure 15).

• For the same reason, the tails develop essentially kinematically, i.e., r = v × t. Although in Chapter IV of this thesis we will illustrate that this is not so for much of the material in galaxies modeled with heavy halos, it remains true along much of the length of the tails.^{3}

• Particles in a prograde disk in a bound orbit will experience the strongest tidal response, due to the “near-resonance or matching of their orbital speeds with the peak angular motion of the companion”. Unbound encounters will have no impetus to match these speeds. Particles in a disk undergoing a retrograde encounter will feel the perturbing potential pull them alternately inward and outward over one orbit, smoothing out the tidal field and creating a much milder response.

• Both tails and bridges form only after orbital periapse. Bridges form from material which streams from one body towards (but not through) the location of the other. They will therefore be destroyed shortly after the galaxies return for a second encounter. Tails, on the other hand, expand into the far side of the potential, which remains relatively quiescent. For this reason, mergers will exhibit one tail per disk and a spray of former bridge material opposite this tail.

• Galaxy mass ratios near one and small disk inclinations make better tails and worse bridges.

3 TT differentiated between bound and unbound antitidal material by the terms counterarm and tail, respectively. I will always use the term tail, as this distinction cannot be made without the aide of a detailed model of the interaction.

2. Historical Perspectives 7

This Page Left For Fig. I-3

FIGURE 4: The Toomre Sequence of eleven NGC prospects for ongoing mergers, as sketched by Toomre (1977). The NGC number of each system is listed to the left, while the columns to the right give the VV and Arp designation (Vorontsov-Vel’yaminov 1959, and Arp 1966, respectively), the right ascension and declination, and the heliocentric velocity (km s^{−1}). Four of the systems chosen for this study were taken from this sequence (NGC 4676, NGC 520, NGC 3921, and NGC

Having established that tails and bridges can result from tidal perturbations, the Toomres (1972) considered what might have happened had their massless test particles actually had mass, and had therefore been capable of transferring energy and angular momentum from the orbiting galaxies to the tidal features, pondering whether perhaps the two galaxies might not “lose altogether their separate identities and simply blend or tumble into a single three dimensional pile of stars” (TT p. 662). This led them to propose several systems as plausible examples of such coalescing objects.

With the aid of deep blue (IIaO and IIIaJ) and near infrared (N) plates taken by F. Schweizer while at the Hale Observatories, Alar Toomre and Schweizer noted the occurrence of multiple nuclei of various separations as well as single nuclei among the multi-armed systems. This led them to refine the list of candidate mergers and to arrange the systems in a suggestive time sequence, from separate strongly interacting systems likely to merge, to systems in the throes of merging, to well merged systems. Toomre’s (1977) table is reproduced here as figure 4, as it is the focal point for this thesis. It represents the foremost 11 examples of ongoing mergers from the NGC, and I will refer to it throughout this thesis as the Toomre Sequence.

From these eleven galaxies out of the 4000+ NGC entries, and with an age estimate for tidal features of 5x10^{8} yrs based on rotational periods, Toomre (1977) calculated that there must be several hundred former mergers among the galaxies in the NGC, a number surprisingly close to the 500 or so ellipticals that are in that catalogue. Based on the hypothesized morphology of the post-merging star-piles and the above statistics, he concluded that elliptical galaxies must be these remnants. This train of thought — that pre-formed pairs of disk galaxies can loose their mutual orbital energy and angular momentum, losing also their separate identities in the process, leaving behind a full fledged elliptical in their stead — is what is now referred to as the Toomre Merger Hypothesis or the Merger Hypothesis.

2.2 Observational Validation

It was the detailed observational studies of apparently interacting or merged systems that added strong support to the merger hypothesis and generated widespread enthusiasm for the field. This work was carried out on several fronts, and I here enumerate those with the most relevance to the present work.

• Detailed studies of individual merger candidates by Schweizer (1978, 1980, 1981, 1982) put the study of mergers on sound observational ground by systematically defining the expected characteristics of mergers and merger remnants and looking for them. This effort is described in more detail in §2.2.1 below.

• Improved technology and careful observations revealed that ellipticals were not as regular as formerly believed (Malin 1977, 1978, 1979; Malin & Carter 1980, 1983), and compelling evidence for both minor and major accretion events has been amassed (cf. Athanassoula & Bosma 1985). This evidence leads to the conclusion that perhaps the majority of ellipticals have had such events in their past. This work is summarized in §2.2.2.

• An influential study on the global colors of peculiar galaxies coupled with simple theoretical models by Larson and Tinsley (1978) led to a simple framework in which to understand the large spread of colors in interacting systems and to the conclusion that such systems had undergone a recent burst of star formation. This led to a growing appreciation that the lives of galaxies are not as quiescent as formerly believed. It is fair to say that this paper had the same stimulating effect on the observational community that the Toomres had upon the theoretical world (see Kennicutt 1990 for review). This was because their model — normal stars, an assumed IMF, and star formation — was simple, explained the observations, and suggested many tests, such as looking at other tracers of star formation and burst strength. The subsequent studies served mainly to verify Larson & Tinsleys’ original conclusion: that interactions can have a drastic effect on the star formation history of the participants.

• A link between induced central activity and galaxy interactions was suggested as early as 1954 for the radio-active galaxy Cen A (Baade & Minkowski 1954). This connection was put on sound statistical grounds by studies which showed a clear increase in the incidence of Syfert nuclei (Adams 1977) and radio activity (Stocke 1978) in galaxies with companions. The studies that this work stimulated eventually led to the conclusion that an external property such as environment could have a deciding role on a galaxies internal properties (see review by Heckman 1990), causing otherwise preoccupied astronomers to pay attention to such supposedly “rare” events. Further, the association of radio, Seyfert, and QSO activity with the presence of tidal distortions, companions, and/or redshifted H i absorption suggests interaction induced gaseous infall as a likely meal for lurking central monsters (see reviews by Heckman 1990, Stockton 1990).

• The 10μm luminosities of suspected mergers were found to be much higher than their optical luminosities (Lonsdale et al. 1984, Joseph & Wright 1985). Soon after, follow-up optical imaging of the brightest extragalactic sources uncovered by the IRAS survey revealed that the vast majority of them were optically very disturbed or strongly interacting (Houck et al. 1985, Cutri & McAlary 1985, Sanders & Mirabel 1985, Sanders et al. 1986, 1988). These systems have bolometric luminosities similar to QSOs (Soifer et al. 1984), leading to speculation that such activity may be triggered gravitationally, either by an induced starburst (Lonsdale et al. 1984) or by feeding a central “monster” (Toomre & Toomre 1972, Gunn 1979). This raises the possibility that nearby mergers are providing a local glimpse of the earliest phase of galactic evolution.

It is the first two and the last of these efforts which made the most direct contributions to the Merger Hypothesis, and I will discuss them in more detail below. The importance of the other two cannot be denied nor underestimated. However, their contribution is less directly related to mergers than to interactions in general. I will therefore not expand on these themes other than the comments made above, referring the interested reader to the listed reviews.

2.2.1 Definition of a Merger Remnant

To put the merger hypothesis on firm ground, F. Schweizer sought to define the observational characteristics of an object that was “in the act” of becoming an elliptical (Schweizer 1978). He knew that the system should have a single nucleus, yet two symmetric tidal features moving in opposite directions. It should be well isolated to remove the possibility that an interloper caused the damage. Finally, from his spectroscopic studies of interacting and merging systems he expected to find chaotic internal motions in the central area. Though most of the galaxies in the Toomre Sequence exhibited many of these qualities, it was NGC 7252, the “Atoms for Peace” galaxy, which possessed them all, and which Schweizer therefore proposed as the prototypical elliptical-in-formation.

Schweizer followed up his claim with a detailed spectroscopic and photometric study of this system (Schweizer 1982). Perhaps the most impressive and important result was that the light profile derived from concentric aperture photometry of this messy system was very well represented by an r^{1/4} law, despite the many shells, loops and ripples seen in Fig. 5b (from Schweizer 1990). Also important was the discovery of a significant Balmer absorption component to the galactic light (the so-called post-burst or E+A type spectra, reproduced here in Fig. 5c, also from Schweizer 1990) which pointed to the occurrence of a central burst of star formation ∼ 10^{8} years previous. Subsequent modeling of this spectrum led to an estimate that between 10–50% of the present stellar content was created in this burst (Fritze von-Alvensleben & Gerhard 1994). NGC 7252 also had one of the kinematically distinct subsystems in its ionized gas, exhibiting both minor axis rotation and major axis velocity reversals. The kinematics along the major and minor axis are shown in Fig. 5a, reproduced from Schweizer (1982). The ionized gas disk has normal rotation out to 20′′, at which point the rotation curve switches rotational sense along the major axis, and exhibits apparent minor axis rotation at the same point. This further strengthened the claim that NGC 7252 was the result of distinct components merging into a single object.

These observations add strong support for the hypothesis that mergers of spirals can make ellipticals. The r^{1/4} or de Vaucouleurs profile (de Vaucouleurs 1953) is the characteristic profile of ellipticals, and thought to be an indication that violent relaxation has occurred (Lynden-Bell 1967). Since the two tails point to two nearly equal-mass disk progenitors, this observation is perhaps the single strongest support for the merger hypothesis. The study presented below in Chapter III reinforces this result by showing that NGC 7252 falls upon the more restrictive fundamental plane relationship for ellipticals (Djorgovski & Davis 1987; see also

FIGURE 5: Observations of NGC 7252: (a) Rotation curves of inner ionized gas disk as measured from CTIO image tube spectrographs, taken from Schweizer (1982). The top panel gives the velocity along the minor axis, which exhibits strong motions at about ±15′′, while the bottom panel presents the major axis rotation curve, which reverses its sense of motion at about ±10′′. (b) Unsharp masked image of the remnant body, showing the many surrounding ripples, loops and shells, taken from Schweizer (1990). (c) Blue spectrum of the remnant nucleus (Nuc) and 14′′ off-nucleus, taken from Schweizer (1990). Both spectra show the Balmer absorption lines characteristic of an intermediate aged stellar population, evidence for a past burst of star formation.

Appendix VI).

The existence of a post-starburst population became even more significant after the discovery of a concentrated disk of molecular gas by Dupraz et al. (1990) and Wang et al. (1992). The mass density of molecular gas in the inner region is comparable to the stellar mass density in ellipticals (Wang et al. 1992), and much higher than found in normal spirals (Kormendy & Sanders 1992). It is evidence that significant dissipation has taken place in the gaseous component of the progenitors (TT, Carlberg 1986, Negroponte & White 1983), and subsequent star formation at the rate inferred by Fritze von-Alvensleben & Gerhard would greatly increase the phase space density of the remnant. This removes one of the oldest objections to the merger hypothesis (Ostriker 1980, Carlberg 1986).

Having successfully defined and tracked down a merger remnant, Schweizer and collaborators then turned their efforts towards finding the objects between NGC 7252 and ellipticals; that is, of finding the very old merger remnants. The existence of such objects is central to the merger hypothesis, as it makes the difference between mergers producing peculiar galaxies and mergers producing ellipticals. These efforts are described in the next section.

2.2.2 Shells, Fine Structure, and Aging Merger Remnants

At the discussion following Toomre’s presentation of his merging sequence, Ivan King inquired about the objects between late stage mergers such as NGC 3921 and NGC 7252 and ellipticals. Surely the sky must teem with these objects, if they are the precursors to all ellipticals, so where are they and what do they look like? The observational work to uncover the occupants of this so called “King Gap” was again led by F. Schweizer. This involved taking the next step beyond the previously defined merger object, and asking what it would become after signs of tidal interaction had faded into the background (Schweizer 1983).

Some clues had emerged with the discovery of faint shells or ripples in roughly one-fourth of field ellipticals by Malin & Carter (1980, 1983). Such faint structures embedded within the bright light and rapidly declining light profiles of the underlying galaxy became observable with the development of “photographic dodging” and “unsharp masking” techniques (Malin 1977, 1978, 1979, Schweizer & Ford 1984). Motivated by observations of similar structures in the well documented merger candidates NGC 1316 and NGC 3921 (Schweizer 1980, 1981, 1982), Schweizer came to suspect that ellipticals themselves may bear the made-by-Toomre label, by exhibiting many examples of merger or merger-induced accretion.

In collaboration with P. Seitzer, Schweizer conducted several surveys of E’s, SO’s, and spirals to look for shells and other “fine structures” (Schweizer & Seitzer 1988, 1992; Seitzer & Schweizer 1990). They found that about half of the ellipticals and one-third of the SO’s exhibit such structure. However, Schweizer’s early work notwithstanding, the connection between fine structure and major mergers was not yet widely accepted. Early numerical work had demonstrated in principle that such features could be produced by small accretion events (Quinn 1984, Dupraz & Combes 1986, Hernquist & Quinn 1988, 1989) and it wasn’t realized until much later that the exact same physics should apply to the cold material falling back in from the tails during major mergers (Hernquist & Spergel 1992). It was therefore necessary to demonstrate a link between the presence of shells and other fine structure features and intrinsic properties of the galaxies which would suggest a not so distant merger past.

To this end, Schweizer et al. (1990) defined a fine-structure index, Σ, to characterize the number of induced features, such as ripples, tails, plumes, X-structures, and boxy isophotes, a galaxy possessed. They then looked at many intrinsic properties of their program ellipticals as a function of Σ. This intuition paid off, as Σ was found to correlate with signs of global youth in the parent system, such as stellar absorption line strengths (Schweizer et al. 1990) and blue UBV colors (Schweizer & Seitzer 1992). This correlation is in the sense that younger looking systems have a higher value of Σ. Knowledge of this correlation allowed a phenomenological age to be derived from burst models and the global colors of the galaxy (Schweizer & Seitzer 1992). From this work came the suggestion of several candidates for residence in the “King Gap” between the Toomre ongoing mergers and ellipticals.

These concepts allow us to judge the steps lying between the present sequence and ellipticals, if indeed such an evolutionary tract exists. In turn, more detailed

FIGURE 6: The merging sample of Joseph & Wright (1985), reproduced from their Plate 1. These systems were selected for their morphological peculiarities, and were subsequently found to be ultraluminous at 10μm. They are arranged according to a putative merging timeline, from bottom to top and left to right. In this order they are (a) NGC 520, (b) NGC 2623, (c) NGC 3256, (d) NGC 1614, (e) IC 883, (f) NGC 6240, (g) Arp 220, (h) NGC 4194, (i) NGC 3310.

observations of youngish remnants such as NGC 3921 and NGC 7252 can serve as a guide to calibrate better these heuristic age estimators.

2.2.3 Ultraluminous IRAS Galaxies and Mergers

The IRAS mini-survey (Neugebauer et al. 1984; Rowan-Robinson et al. 1984) revealed a class of objects which emitted well over 100 times as much in the infrared than in the visible (Soifer et al. 1984, Houck et al. 1985), making them 100-1,000 times brighter than normal spirals (de Jong et al. 1984). The optical identification of these objects revealed that the majority were peculiar (Houck et al. 1985, Sanders et al. 1986) and, based upon their molecular gas content and optical morphologies that included wisps, tails, jets, rings, and/or double nuclei, it was suggested that the most extreme infrared emitters were ongoing mergers of gas-rich systems (Sanders et al. 1986, 1988). These ultraluminous infrared (ULIR) galaxies are characterized by far-infrared luminosities of L_{IR}_{∼}^{>} 3×10^{11} h^{−2} L_{⊙}^{4}.

At around the same time, Joseph & Wright (1985) investigated the IR properties of a restricted merging sample. This sample was composed of all equal mass disk-disk mergers in which the two disks have already coalesced into a single objects for which 10μm luminosities were available. This sample of nine galaxies is reproduced here in figure 6. The left-most three systems (Fig. 6a,b,c) are numbers 7–9 from the Toomre Sequence (Fig. 4). All nine of the candidate mergers were found to be ultraluminous at 10μm (L_{10μm} > 10^{9} h^{−2} L_{⊙}; for comparison the 10μm luminosities for bright spirals is given as 0.3–2×10^{8} h^{−2} L_{⊙} and for Seyferts as 1– 300×10^{8} h^{−2} L_{⊙}). From the IR colors and continuum spectra of the mergers and the lack of evidence for a dominant central engine, these authors concluded that the IR luminosity was being powered by a merger induced super starburst (see also Lonsdale et al. 1984). They further suggested that this super starburst would impart sufficient energy into the remaining ISM to sweep the remnant clean (Graham et al. 1984), leaving objects which would “indeed become indistinguishable from elliptical galaxies” (Joseph & Wright 1985, p. 93).

Given the proposed link between the ULIR systems and gas rich mergers, Joseph & Wright (1985) proposed that the optical morphology and IR luminosity could be used place systems within the following evolutionary scenario: gasrich galaxies collide, a starburst ensues, the IR luminosity increases, the starburst consumes the central gas supply and dies down, and the IR luminosities return to normal levels. This scenario was illustrated graphically by Casoli et al. 1991 (their figure 9), reproduced here as figure 7. This figure charts the behavior of several measures of IR activity along a merging timeline similar to that described above: from spirals, through five systems from the sample of Joseph & Wright (NGC 3256=Fig. 6c, NGC 6240=Fig. 6f, NGC 1614=Fig. 6d, Arp 220=Fig. 6g), to a merger remnant (NGC 7252), and finally to an elliptical with a

4 As calculated for the 42.5–122.5 μm band, from the 60μm and 100μm IRAS fluxes using the prescription given in Helou et al. (1988). Note: Due to different adopted distance scales and methods of calculating L_{IR}, the absolute luminosity corresponding to “ultraluminous” varies according to author.

FIGURE 7: A graphic illustration of the behavior of the IR properties along the suggested merger timeline of Joseph & Wright (1985), taken from figure 9 of Casoli et al. (1991). The measures of IR activity are described in the text. According to the ULIR-merger paradigm, two gas-rich systems fall together, stimulating a central starburst. The massive stars formed in this burst heat the surrounding dust, which radiates this energy away efficiently at IR wavelengths. As a result, the IR luminosity is increased to 10–100 times the pre-encounter levels. In this example, NGC 3256 is just beginning its IR luminous phase, NGC 1614 and Arp 220 are at the peak of their starburst, the burst in NGC 2623 is just past the peak, while NGC 7252 is in a post-burst stage (appropriate to its nuclear spectrum given in Fig. 5). The final system, NGC 3928, is an elliptical galaxy with a large amount of fine structure. While the horizontal axis is meant to represent a time-ordered sequence, time is not meant to increase linearly along this axis, i.e. the time for a system to evolve from the NGC 1614 stage to the Arp 220 stage is not meant to be equal to the time to evolve from the NGC 7252 stage to the NGC 3928 stage.

FIGURE 8: An alternative evolution of the IR luminosity to the timeline of Casoli et al. (1991) given in Fig. 7. This figure present the results of a numerical simulation between equal-mass disks, using a formulation for gas dynamics and star formation, taken from Mihos & Hernquist (1994, their figure 2). The star formation rate (∼IR luminosity) as a function of time is plotted for two encounter which differed only by the structure of the progenitors. The thin line gives the burst evolution of a merger of bulgeless-disks. First pericenter occurs at about T=32, stimulating rapid inflow of gas and stimulating a mild burst. This burst consumes most of the gas of the progenitors, so there is little to fuel a burst at the final coalescence of the nuclei (T≈70). The dark line shows the burst evolution of a merger of disk galaxies with a bulge. In this case, the bulge inhibits gas inflow to the center, and as a result the disks remain gas-rich until the final merger at T≈65, fueling a greatly increased burst of star formation. This simulation cautions us that there may be necessary pre-conditions on the progenitors in order for a merger to lead to an ULIR phase. This idea is addressed in Chapter VI of this thesis.

large amount of fine structure (NGC 3928). The measures of IR activity are the farinfrared luminosity per unit blue luminosity (L_{IR}/L_{B}), believed to be a measure of the massive star formation rate (Solomon & Sage 1988), the IR luminosity per unit molecular gas mass (L_{IR}/M_{H_{2}}), also referred to as the “star formation efficiency” as it is a measure of the number of massive stars being formed per molecular cloud (Young et al. 1986, Sanders et al. 1986, Solomon & Sage 1988), and the far-IR color temperature (T_{60/100}), reflecting the temperature of the starburst-heated dust.

It is important to realize that this timeline is purely heurestic, and it is quite possible that the systems plotted in Fig. 7 are each on very different evolutionary paths. It may well be that only certain types of progenitors or encounter geometries will lead to an ULIR phase. To illustrate this point, we reproduce in figure 8 a plot of the star formation history of two simulated mergers, taken from figure 2 of Mihos & Hernquist (1994). These simulations included a prescription for both gas dynamics and star formation, and the two runs were identical except for the internal structure of the progenitors. These authors found that progenitors without a bulge would experience a much milder and longer lived burst of star formation (the Disk/Halo line in Fig. 8) than systems with a bulge (the Disk/Bulge/Halo line in Fig. 8). If this case is indicative, Fig. 7 may be the superposition of many parallel sequences at different stages of evolution, with e.g. NGC 7252 at the tail of the Disk/Halo (i.e. no bulge) evolutionary path in Fig. 8, and Arp 220 at the peak of Disk/Bulge/Halo evolutionary path. This is an important issue to resolve if we are to understand the relationship between disk-disk mergers such as the Toomre Sequence and the ULIR galaxies.

3. Formulation of The Thesis

The Merger Hypothesis had been advanced in 1972 and more formally in 1977 (Toomre & Toomre 1972, Toomre 1977). And yet, as of 1988 when I came to Columbia University, the only major merger from the Toomre Sequence which had been mapped in neutral hydrogen was NGC 4038/9, “The Antennae” (van der Hulst 1979). This system has a gas rich tail leading south to a dwarf irregular-like clump at its tip (cf. Schweizer 1978), a gas poor tail leading to the north, and a significant amount of atomic gas within both highly distorted disks (see also Mahoney et al. in preparation). If this system is to evolve into an atomic gas poor elliptical, then the ultimate fate of its atomic gas has to be addressed

But had Schweizer (1982) so convinced the community that mergers would structurally make ellipticals from spirals, and had Toomre & Toomre convinced us that the most gas rich regions would be forever ejected into freely expanding tidal features, and Joseph & Wright (1985) convinced us that any central gas would either be converted into stars or swept from the remnant by the ensuing starburst, that observational evidence to this end was merely a formality? At the Heidelberg meeting on the “Dynamics and Interactions of Galaxies” (Wielen, 1990), François Schweizer, Jacqueline van Gorkom and I decided that the answer to this question was no, and further decided that systems drawn from the Toomre Sequence would make an ideal sample for a systematic study to determine experimentally the fate of the neutral hydrogen at progressively advanced stages of merging.

Recent hydrodynamical simulations suggest that most of the gas will very quickly accumulate in the center (Schmutzler and Biermann 1985, Hernquist 1989, Noguchi 1988, Barnes & Hernquist 1991). This seems consistent with the presence of large amounts of molecular gas in the centers of bright IRAS galaxies (Scoville et al. 1991), which are thought to be merger products (§2.2.3). On the other hand, in interacting galaxies the neutral hydrogen is often very extended, stretching over several optical galactic diameters (van der Hulst 1979, Haynes et al. 1984, Simkin et al. 1987, Appleton et al. 1987, Rots et al. 1990, Yun et al. 1994), providing more evidence for gas being pulled out than falling in. What then happens to the gas in major mergers, where optically we see such long, narrow ribbons of starlight stretching for tens of kiloparsecs into intergalactic space? Since tails are drawn from the outer regions of disk galaxies (Fig. 3), and these regions tend to be quite rich in neutral hydrogen (Wevers et al. 1986), mapping in the 21cm line of neutral

§3. Formulation of The Thesis 19

hydrogen allows us to measure both the distribution and kinematics of the atomic gas in such systems.

For these reasons it was decided to use the spectral line mapping capability of the 27 element radio synthesis telescope of the National Radio Astronomical Observatory (NRAO), the Very Large Array (VLA; Napier et al. 1983). Relevant observational details of this instrument are described in Appendix I at the end of this thesis. By combining data from two configurations we are able to achieve spatial resolutions of order 15′′ and column density sensitivities of order 10^{19} cm^{−2}. This enables us to not only measure the quantities of central atomic gas and to completely map of the kinematics of the tidal regions, but also to survey a large volumes of space (∼0.5 Mpc in diameter) surrounding each system. This is important if we are to have faith that we have detected all of the remaining atomic gas from the progenitors.

We are also interested in process occurring within the tidal features, particularly in any that differentiate between gas and stars. In NGC 4038/9, the morphology of the tails shows gaps, clumps and a possibly embedded dwarf. The difference between the H i distribution in the northern and southern tails is quite remarkable, given their similar optical appearance, and suggest that forces other than gravity are at work. Additionally, giant knots of star formation are frequently found within tidal tails (Schweizer 1978), and tails and bridges as a class have optical colors indicating in situ star formation (Schombert et al. 1990). It was therefore decided to use narrow band imaging in Hα and wide-field broad-band imaging in BV R to investigate what role, if any, star formation plays within the tails. Hα will locate sites of current star formation, while the BV R colors will identify intermediate aged populations. Wide-field imaging allows the underlying colors to be correlated with the local H i content and kinematics. This combination of data — neutral hydrogen kinematical mapping, narrow-band Hα and wide-field broad-band BV R imaging, should be well suited to evaluate the importance of such processes as self-gravity and dissipation on the dynamics of the gas within the tails.

Originally this thesis was to be kept purely observational. However, drastic improvements in the ability to run numerical simulations over the short timespan from when I first entered Columbia University in 1988 until the year before my defense (1993) enabled me to realize a personal dream: to constrain detailed numerical models of a merging encounter using an H i data cube. These improvements came on many fronts (cf. Barnes & Hernquist 1992 and references therein). On the software front, algorithms utilizing tree structures that were developed in the mid-80s (e.g. Barnes & Hut 1984) were refined to the point that such codes could be run by non-specialist. These algorithms reduce the required number of force calculations by an order of magnitude, allowing self-consistent calculations including multi-component galaxies and hydrodynamics to be run in reasonable times. Additionally, much of the dirty work in learning how to build “realistic” numerical galaxies and investigating the stability of such systems had already been done (e.g. Barnes & Hut 1989, Hernquist 1993, and references therein).

On the hardware front, the size and cost of both the processors and the storage devices needed to conduct such numerical experiments dropped precipitously, to the point that particularly generous advisors could afford to provide their graduate students with hardware that would occupy a fraction of a desk-top and could run a simulation overnight. All that was missing was software enabling a direct comparison between the three-dimensional H i data and the six-dimensional simulation snapshots. Another goal of this thesis is to develop such a tool.

Even though we had enlisted the help of several experts for our foray into realms theoretical (namely L. Hernquist and J.C. Mihos of UCSC), modeling merging encounters is still a daunting undertaking. There are a dizzying number of free parameters, and the dynamical effects in the inner regions — including gas viscosity, cloud collisions, shocks, and star formation — are poorly understood. Attempts to model the inner regions of interacting systems have proven problematic (cf. Hernquist 1990). However, restricting our model matching to the kinematics of the distant tidal regions offers more promise for the following reasons: the general potential should be much better behaved than in the rapidly evolving central regions; the areas are very well delineated; the timescales can be well constrained from kinematic data alone; non-linear effects should be minimized since the velocity fields and dispersions are much more ordered; and the tidal material is passing through the region for presumably the first time.

The aim of this thesis, then, is to obtain high quality radio and optical data for a proposed evolutionary sequence of disk-disk mergers. These data will be used to determine observationally the effects of the interaction on the gaseous component of the progenitors, and to constrain numerical models of the interaction. These models will help us explore the ultimate fate of the gas during the merging process.

Since this is a two-pronged investigation, there are two types of questions I wish to address in this thesis.

Observationally, I will ask:

(1) Do these systems indeed appear to be spirals or former spirals colliding? If so, does it seem likely that the end product will be an elliptical?

(2) Do the chosen system in fact form sequence? If so, are there any trends along the sequence?

(3) Are there any differences between the gas and stars within these systems? Can we understand them?

(4) What are the substructures within the tails? Can we understand why they are there and/or how they will evolve?

Dynamically, I wish to know:

(5) Is Toomre & Toomre 1972 the whole story?

(6) Can the tail kinematics usefully constrain the encounter geometry? If so,

§4. Layout of the Thesis 21

what can we learn about both the past and future of the encounter, especially as concerns the atomic gas?

(7) Does the combination of simulations and data give us any new information that was not already apparent in either one alone?

(8) What can the observed trends and the dynamics of this sequence tell us about the connection between merging and the ULIR galaxies?

4. Layout of the Thesis

The systems chosen for this study are presented in Chapter II. The observational data and reduction techniques are then described. In the results section we discuss the data on each object in a cursory manor, with special attention to trends along the sequence, especially as concerns any differences between the gas and starlight. In the last section we discuss these results in terms of the evolutionary sequence and attempt to extrapolate beyond the sequence to predict the fate of the remnants.

In Chapter III, I utilize the critical analysis and writing talents of Puragra Guhathakurta, Jacqueline van Gorkom and François Schweizer to help present and disentangle the wealth of data collected on NGC 7252. We go into depth on all aspects on this merger remnant with a detailed study of the H i, H ii, B, R, and x-ray data of this system. The analysis of the remaining systems in the sequence will be patterned after this study. This study was published in the Astronomical Journal (1994, AJ, 107, 67).

In Chapter IV, I enlist the aide of J. Christopher Mihos to help use the H i kinematical data of NGC 7252 to constrain N-body simulations of the original encounter. Lars Hernquist kindly donated the CPUs for this project. The data are successfully fit, and the evolution is followed both forward and backward in time to better understand the orbital structure of the tails. This information is used to help judge the future evolution of NGC 7252. This chapter was submitted to the Astronomical Journal in December of 1994.

In Chapter V, I will ask whether disk-disk mergers such as the five systems studied here appear to conform to the paradigm illustrated in Fig. 7, or if a situation similar to that illustrated in Fig. 8 is more applicable. That is, I will investigate whether a condition other than just merging is need to fuel an ULIR phase. While the published data is suggestive of such a precondition, I cannot definitively answer this question, and therefore outline several avenues of research which should.

In Chapter VI I summize the results of the previous six chapters and address the eight questions posed in §3 of this chapter. I conclude with a description of what I think are the four most interesting lines of research for future studies.

In the Appendices I present supplementary information on both the observations and simulations which did not make it into Chapters II–V due to space constraints. In the first Appendix, I present added details on the capabilities of the VLA and the strategy employed for the H i observations. In Appendices II–V, I present the H i channel and moment maps for each system except NGC 7252, and more detailed images illustrating some of the idiosyncrasies referred to in Chapter II. In Appendix VI, I show the photometric properties of NGC 7252 compared to ellipticals and spirals, and present B-R color maps of both NGC 7252 and NGC 4038/9, another member of the Toomre Sequence that was observed over the course of this work. In Appendix VII, I present a brief summary of the properties of several clumps found within the tails of the late stage mergers.

Finally, in Appendix VIII, I describe the data matching process used in Chapter IV in more detail, and elaborate on the mechanics of tail formation and very long term evolution.

References

Adams, T.F., 1977, ApJ (Supp.), 33, 19. Alladin, S.M., 1965, ApJ, 141, 768. van Albada, T.S., van Gorkom, J.H, 1977, A&A, 54, 121. Appleton, P.N., Ghigo, F.D, van Gorkom, J.H., Schombert, J.M., Struck-Marcell, C., 1987,

Nature, 330, 140. Arp, H. C., 1966, “Atlas of Peculiar Galaxies” (Pasadena: CalTech). Athanassoula, E., Bosma, A., 1985, ARA&A, 23, 147. Baade, W., Minkowski, R., 1954, ApJ, 119, 206. Barnes, J.E., 1994, in “The Formation of Galaxies: Proceedings of the V Canary Islands Winter

School of Astrophysics”, ed. C. Muñoz–Tuñón, (Cambridge: Cambridge University Press), in press.

Barnes, J.E., Hernquist, L., 1991, ApJ (Lett.), 370, L65. Barnes, J.E., Hernquist, L., 1992, ARA&A, 30, 705. Barnes, J.E., Hut, P., 1986, Nature, 324, 446. Burbidge, E.M., Burbidge, G.R., Hoyle, F., 1963, ApJ, 138, 873. Carlberg, R.G., 1986, ApJ, 310, 593. Casoli, F., Dupraz, C., Combes, F., Kazes, I., 1991, A&A, 251, 1. Cutri, R.M, McAlary,C.W., 1985, ApJ. 296, 90. Djorgovski, S., Davis, M., 1987, ApJ, 313, 59. Dupraz, C., Casoli, F., Combes, F., Kazes, I., 1990, A&A, 228, L5. Dupraz, C., Combes, F., 1986, A&A, 166, 53. Fritze-von Alvensleben, U., Gerhard, O.E., 1994, A&A, 285, 751 & 775. Graham, J.R., Wright, G.S., Meikle, W.P.S., Joseph, R.D., 1984, Nature, 310, 213. Gunn, J.E, 1979, “Active Galactic Nuclei”, eds. C. Hazard and S. Mitton (Cambridge: Cam-

bridge Univ. Press), p. 213. Haynes, M.P., Giovanelli, R., Chincarini, G.L, 1984, ARA&A, 22, 445. Heckman, T.M., 1990, in “Paired and Interacting Galaxies”, IAU Colloq. No. 124, eds. J.W. Su-

lentic, W.C. Keel, and C.M. Telesco (Wash. D.C.: NASA-CP-3098), p. 357. Helou, G., Khan, I.R., Ladan, M., Boehmer, L., 1988, ApJ (Supp.), 68, 151. Hernquist, L., 1989, Nature, 340, 687. Hernquist, L., 1990, in “Dynamics and Interactions of Galaxies”, ed. R. Wielen (Heidelberg:

Springer), p. 108. Hernquist, L., 1993b, ApJ (Supp.), 86, 389. Hernquist, L., Quinn, P.J., 1988, ApJ, 331, 682. Hernquist, L., Quinn, P.J., 1989, ApJ, 342, 1. Hernquist, L., Spergel, D.N., 1992, ApJ (Lett.), 399, L117. Holmberg, E., 1941, ApJ, 94, 385. Houck, J.R., Schneider, D.P., Danielson, G.E., Beichman, C.A., Lonsdale, C.J., Neugebauer, G.,

Soifer, B.T., 1985, ApJ (Lett.), 290, L5. van der Hulst, J.M., 1979, A&A, 155, 151. Joseph, R.D., Wright, G.S., 1985, MNRAS, 214, 87. de Jong, T., Clegg, P.E., Soifer, B.T., Rowan-Robinson, M., Habing, H.J., Houck, J.R., Aumann,

H.H., Raimond, E., 1984, ApJ (Lett.), 278, L67. Kennicutt, R.C. Jr., 1990, in “Paired and Interacting Galaxies”, IAU Colloq. No. 124, eds.

J.W. Sulentic, W.C. Keel, and C.M. Telesco (Wash. D.C.: NASA-CP-3098), p. 269. Kormendy, J. Sanders, D.B., 1992, ApJ (Lett.), 390, L53. Larson, R.B., Tinsley, B.M., 1978 ApJ, 219, 46. Lonsdale, C.J., Persson, S.E., Matthews, K., 1984, ApH, 287, 95. Lynden-Bell, D., 1967, MNRAS, 136, 101. Malin, D. F., 1977, AAS Photo Bull., 16, 10. Malin, D. F., 1978, Nature, 276, 591. Malin, D. F., 1979, Nature, 277, 279. Malin, D. F., Carter, D., 1980, Nature, 285, 643. Malin, D. F., Carter, D., 1983, ApJ, 274, 534. Mihos, J.C., Hernquist, L., 1994, ApJ (Lett.), 431, L9. Napier, P.J., Thompson, A.R., Ekers, R.D., 1983, Proceedings IEEE 71, p. 1295. Negroponte, J., White, S.D.M., 1983, MNRAS, 205, 1009. Neugebauer, G., et al., 1984, ApJ (Lett.), 278, L1. Noguchi, M., 1988, A&A, 201, 37. Ostriker, J.P., 1980, Comments on Astrophysics, 8, 177. Pikel’ner, S.B., 1968, ARA&A, 6, 165. Quinn, P.J. 1984, ApJ, 279, 596. Rots, A.H., Bosma, A., van der Hulst, J.M., Athanassoula, E., Crane, P.C., 1990, AJ, 100, 387. Rowan-Robinson, M., et al., 1984, ApJ (Lett.), 278, L7. Sanders, D.B., Mirabel, I.F., 1985, ApJ (Lett.), 298, L31. Sanders, D.B., Scoville, N.Z., Young, J.S., Soifer, B.T., Schloerb, F.P., Rice, W.L., Danielson,

G.E., 1986, ApJ (Lett.), 305, L45. Sanders, D.B., Soifer, B.T., Elias, J.H., Madore, B.F., Matthews, K., Neugebauer, G., Scoville,

N.Z., 1988 ApJ, 325, 74. Schmutzler, T., Biermann, P., 1985, A&, 149, 383. Schombert, J.M., Wallin, J.F., Struck-Marcell, C., 1990, AJ, 99, 497. Schweizer, F., 1978, “The Structure and Properties of Nearby Galaxies”, IAU Symp. No. 77,

eds. E.M. Berkhuijsen and R. Wielebinski (Dordrecht: Reidel), p. 279. Schweizer, F., 1980, ApJ, 237, 303. Schweizer, F., 1981, ApJ, 246, 722. Schweizer, F., 1982, ApJ, 252, 455. Schweizer, F., 1983, in “The Internal Kinematics and Dynamics of Galaxies”, IAU Symp. No. 100,

ed. E. Athanassola (Dordrecht: Reidel), p. 319. Schweizer, F., 1990. in “Dynamics and Interactions of Galaxies”, ed. R. Wielen (Heidelberg:

Springer), p. 60. Schweizer, F., Ford, W.K., Jr., 1984, “New Aspects of Galaxy Photometry”, Lecture Notes in

Physics No. 232, ed. J.-L. Nieto (Berlin: Springer), p. 145. Schweizer, F., Seitzer, P., 1988, ApJ, 328, 88. Schweizer, F., Seitzer, P., 1992, AJ, 104, 3. Schweizer, F., Seitzer, P., Faber, S.M., Burstein, D., Dalle Ore, C.M., Gonzalez, J.J. 1990, ApJ

(Lett.), 364, L33. Schweizer, F., Seitzer, P., 1992, AJ, 140, 1039. Scoville, N.Z., Sargent, A.I., Sanders, D.B., Soifer, B.T., 1991 ApJ, 366, L5. Seitzer, P., Schweizer, F., 1990, “Dynamics and Interactions of Galaxies”, ed. R. Wielen (Heidel-

berg: Springer), p. 270. Simkin S.M., van Gorkom, J.H., Hibbard, J.E., Hong-Jun, S., 1987, Science, 235, 1367. Soifer, B.T., et al., 1984, ApJ (Lett.), 278, L71. Solomon, P.M., Sage, L.J., 1988, ApJ, 334, 613. Stocke, J.T., 1978, AJ, 83, 348. Toomre, A., 1977, in “The Evolution of Galaxies and Stellar Populations”, eds. B.M. Tinsley

and R.B. Larson (New Haven: Yale Univ.), p. 401. Toomre, A., Toomre, J., 1972, ApJ, 178, 623. de Vaucouleurs, G., 1953. MNRAS, 113, 134. Vorontsov-Vel’yaminov, B.A., 1959, “Catalogue of Interacting Galaxies” (Moscow). Wang, Z., Schweizer, F., Scoville, N.Z., 1992, ApJ, 396, 510. Wevers, B.H.M.R., van der Kruit, P.C., Allen, R.J., 1986, A&AS, 66, 505. White, S.D.M., 1978, MNRAS, 184, 185. Wielen, R., 1990, “Dynamics and Interactions of Galaxies” (Heidelberg: Springer). Young, J.S., Kenney, J.D., Tacconi, L., Claussen, M.J., Huang, Y.-L., Tacconi-Garman, L., Xie,

S., Schloerb, F.P., 1986, ApJ (Lett.), 311, L17. Yun, M.S., Ho, P.T.P., Lo, K.Y., 1994, Nature, 372, 530.

Zwicky, F., 1950, Experientia, 6, 441. Zwicky, F., 1956, Ergebnisse der Exakten Naturwissenschaften, 29, 344. Zwicky, F., 1964, ApJ, 140, 1467. Zwicky, F., 1970, “Advances in Astr. & Astrophy.”, 6, 7, (NYC: Academic Press). Zwicky, F., 1971, “Catalogue of Selected Compact Galaxies and of Post-Eruptive Galaxies”

(Guemligen: Zwicky).